OpenStudy (anonymous):

A die is roll .What is probability of rolling a 5 or a number greater than 3 ?

OpenStudy (anonymous):

rolling a 5 is 1 out of 6 rolling greater than 3 is 3 out of six

OpenStudy (sheetal):

rolling 5 is 1 out of 6 or rolling greater than 3 is 3(4,5,6) out of 6. so req probability=6*5*4*3*2*1/5*4*3*2*1+6*5*4*3*2*1/3*2*1*3*2*1 =6+20=26 ans.

OpenStudy (anonymous):

I'm probably wrong, seems to me that you are going to get a 4,5 or a 6 a half of the time and this includes the 5 case.

OpenStudy (anonymous):

probability of 5 is 1/6 and that of greater than 3 is 1/2